# Local Differential Privacy for Distributed Stochastic Aggregative Optimization with Guaranteed Optimality

### v2

@@ I Introduction
Distributed aggregative optimization addresses cooperative optimization and control problems where each agent’s objective function depends on both its local optimization variable and an aggregate of all agents’ optimization variables. It has found various applications, with typical examples including warehouse placement optimization [1], multi-vehicle charging coordination [2, 3], cooperative robot surveillance [4], and personalized machine learning [5]. To date, plenty of distributed aggregative optimization approaches have been proposed, for both time-invariant [1, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15] and time-varying [4, 5, 16] objective functions. These approaches typically require participating agents to have access to accurate gradients of the objective functions, which, however, may not be feasible in practical applications. For example, in machine learning applications involving massive datasets, gradients are typically computed using randomly sampled data, which inherently introduces noise. Moreover, the use of mini-batch computations adds further randomness, increasing the variability in the gradient estimates available to the agents [17, 18, 19, 20]. These challenges necessitate a reevaluation of traditional approaches that assume accurate gradient information, thereby motivating the development of distributed stochastic aggregative optimization algorithms.
@@
However, developing an effective distributed stochastic algorithm for aggregative optimization is nontrivial due to the unique compositional structure of the objective functions. Specifically, unlike traditional distributed stochastic optimization [17, 18, 19, 20], the objective functions in distributed aggregative optimization explicitly depend on an inner-level aggregative function, which naturally arises in many practical applications. For example, in multi-vehicle charging coordination [2], each user’s cost (objective) function typically relies on electricity pricing, which usually varies as a function of the aggregated charging demand. This compositional dependence makes the gradient computation of objective functions inseparable from the estimation of the aggregative function. In the presence of noise, this inseparability leads to tangled errors in computed gradients and aggregative function estimates, substantially complicating the analysis. Moreover, in practice, computations must be conducted on sampled data, which introduces additional complexity. This is because at each iteration, both the aggregative function and the objective functions are defined by the sampled data, making them inherently time-varying and interdependent across iterations. Consequently, this creates significant challenges for algorithm design and convergence analysis. Although some progress has been made to address noise in distributed aggregative optimization [21, 22], these results only consider the special case where the objective function is stochastic and ignores noise in aggregative function estimation, which significantly simplifies the convergence analysis. In fact, when there is noise in aggregative function estimation, the existing approach in [21] will be subject to steady-state optimization errors. To the best of our knowledge, no results are available that can address stochasticity in both the objective function and the aggregative function in distributed aggregative optimization.
@@
[…]
@@ II-A Distributed stochastic aggregative optimization
We consider $m$ agents cooperatively solving the following aggregative optimization problem:
$\min_{\boldsymbol{x}\in\Omega}~{}F(\boldsymbol{x})\!=\!\sum_{i=1}^{m}f_{i}(x_{i},g(\boldsymbol{x})),~{}~{}g(\boldsymbol{x})\!=\!\frac{1}{m}\sum_{i=1}^{m}g_{i}(x_{i}),$ (1)
where $\boldsymbol{x}=\mathrm{col}(x_{1},\cdots,x_{m})$ denotes the global optimization variable consisting of $x_{i}\in\Omega_{i}$ with $\Omega_{i}\subseteq\mathbb{R}^{n_{i}}$ denoting the constraint set, $g(\boldsymbol{x}):\Omega\mapsto\mathbb{R}^{r}$ is an aggregative function with $\Omega$ denoting the Cartesian product of all $\Omega_{i}$ (i.e., $\Omega=\Omega_{1}\times\cdots\times\Omega_{m}$) and each $g_{i}(x_{i}):\Omega_{i}\mapsto\mathbb{R}^{r}$ being the agent $i$’s private function, and $f_{i}(x_{i},g(\boldsymbol{x})):\Omega_{i}\times\mathbb{R}^{r}\mapsto\mathbb{R}$ is the local objective function of agent $i$. In machine learning applications, the functions $f_{i}$ and $g_{i}$ are typically given by
@@
Assumption 1 is commonly used in distributed aggregative optimization [21, 4, 16] and nonconvex optimization with a compositional structure [54, 55, 56]. Unlike existing distributed aggregative optimization results, in, e.g., [5, 4, 1, 8, 9, 11, 12, 14, 15, 6, 10, 16, 13, 21, 22] that require the objective functions to be strongly convex or convex, we allow both $f_{i}$ and $g_{i}$ to be nonconvex.
@@
$W\!=\!\{w_{ij}\}\!\in\!\mathbb{R}^{m\times m}$ satisfies $\mathbf{1}_{m}^{\top}W\!=\!\mathbf{0}_{m}^{\top}$, $W\mathbf{1}_{m}\!=\!\mathbf{0}_{m}$, and $\|I_{m}+W-\frac{\boldsymbol{1}_{m}\boldsymbol{1}_{m}^{\top}}{m}\|<1.$
Assumption 3 ensures that the graph induced by the matrix $W$ is connected, i.e., there is a path in the graph from each agent to every other agent. Moreover, according to Lemma 4 in [40], we have $\|I_{m}+W-\frac{\boldsymbol{1}_{m}\boldsymbol{1}_{m}^{\top}}{m}\|<1-|\rho_{2}|$, where $\rho_{2}$ is $W$’s second largest eigenvalue.
@@ II-B Local differential privacy (LDP)
$\vspace{-0.2em}\mathbb{P}[\mathcal{A}_{i}(\mathcal{D}_{i},\theta_{-i})\in\mathcal{O}_{i}]\leq e^{\epsilon_{i}}\mathbb{P}[\mathcal{A}_{i}(\mathcal{D}_{i}^{\prime},\theta_{-i})\in\mathcal{O}_{i}].$
@@ III Locally Differentially Private Distributed Aggregative Optimization Algorithm
We propose Algorithm 1 to solve the stochastic aggregative optimization problem (1) while ensuring rigorous $\epsilon_{i}$-LDP with a finite cumulative privacy budget in the infinite time horizon. The injected DP noises satisfy the following assumption:
@@
$\max_{i\in[m]}\{\varsigma_{i,x}\}<v_{x}-v_{z},~{}\max_{i\in[m]}\{\varsigma_{i,y}\}<v_{y},~{}\text{and}~{}\max_{i\in[m]}\{\varsigma_{i,z}\}<v_{z},\vspace{-0.3em}$
@@
with $v_{x},v_{y},v_{z}\in(0,1)$ being the decaying rates of the stepsizes $\lambda_{x}^{t}$, $\lambda_{y}^{t}$ and $\lambda_{z}^{t}$, respectively, in Algorithm 1.
@@
In line 6 of Algorithm 1, we use dynamic average-consensus-based gradient descent to enable each agent $i$ to estimate the optimization variables of all other agents, i.e., $\boldsymbol{x}_{i}^{t}\!=\!\mathrm{col}\{x_{i(1)}^{t},\cdots,x_{i}^{t},\cdots,x_{i(m)}^{t}\}$, where $x_{i}^{t}$ denotes agent $i$’s optimization variable and $x_{i(j)}^{t}$ denotes its estimate of agent $j$’s optimization variable for all $j\!\neq\!i$ and $j\!\in\![m]$. This estimation of neighbors’ optimization variables is an important novelty in our design (absent in all existing distributed aggregative optimization algorithms [1, 5, 4, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 21, 22, 16]), and it is key to ensuring a finite cumulative privacy budget in the infinite time horizon. Specifically, by employing average-consensus on $\boldsymbol{x}_{i}^{t}$, we can ensure a diminishing algorithmic sensitivity (see Eq. (15)), which, combined with meticulously designed diminishing DP-noise variances, can ensure a finite cumulative privacy budget in the infinite time horizon (see Eq. (20)). In contrast, in existing algorithms, the difference $\|x_{i}^{t}-x_{i}^{\prime t}\|$ of state trajectories under adjacent datasets will grow over time, which will in turn lead to a growing sensitivity, and hence, an exploding cumulative privacy budget as iterations proceed.
1:Input: Random initialization $\boldsymbol{x}_{i}^{0}\in\Omega$, $y_{i}^{0}\in{\mathbb{R}^{r}}$, and $z_{i}^{0}\in{\mathbb{R}^{r}}$; matrix $R_{i}\!=\![\boldsymbol{0}_{n_{i}\times n_{1}},\cdots,I_{n_{i}},\cdots,\boldsymbol{0}_{n_{i}\times n_{m}}]\!\in\!\mathbb{R}^{n_{i}\times n}$; stepsizes $\lambda_{x}^{t}=\frac{\lambda_{x}^{0}}{(t+1)^{v_{x}}}$, $\lambda_{y}^{t}=\frac{\lambda_{y}^{0}}{(t+1)^{v_{y}}}$, and $\lambda_{z}^{t}=\frac{\lambda_{z}^{0}}{(t+1)^{v_{z}}}$ with $\lambda_{x},\lambda_{y},\lambda_{z}>0$ and $v_{x},v_{y},v_{z}\in(0,1)$; DP noises $\chi_{i}^{t}\in\mathbb{R}^{r}$, $\zeta_{i}^{t}\in\mathbb{R}^{r}$, and $\boldsymbol{\vartheta}_{i}^{t}\in\mathbb{R}^{n}$ (with $n=\sum_{i=1}^{m}n_{i}$; here, we use bold font to represent stacked vectors of $m$ agents) satisfying Assumption 4.
2:for $t=0,1,\cdots,T-1$ do
4:     $y_{i}^{t+1}=y_{i}^{t}+\sum_{j\in\mathcal{N}_{i}}w_{ij}(y_{j}^{t}+\chi_{j}^{t}-y_{i}^{t})+\lambda_{y}^{t}g_{i}^{t}(R_{i}\boldsymbol{x}_{i}^{t})$;
6:             $+\lambda_{z}^{t}\nabla_{y}f_{i}^{t}(R_{i}\boldsymbol{x}_{i}^{t},\tilde{y}_{i}^{t})$, with $\tilde{y}_{i}^{t}\triangleq\textstyle\frac{1}{\lambda_{y}^{t}}(y_{i}^{t+1}-y_{i}^{t})$;
7:     $\boldsymbol{x}_{i}^{t+1}=\Pi_{\Omega}\Big{\{}\boldsymbol{x}_{i}^{t}+\sum_{j\in\mathcal{N}_{i}}w_{ij}(\boldsymbol{x}_{j}^{t}+\boldsymbol{\vartheta}_{j}^{t}-\boldsymbol{x}_{i}^{t})$
8:             $-\lambda_{x}^{t}R_{i}^{\top}\big{(}\nabla_{x}f_{i}^{t}(R_{i}\boldsymbol{x}_{i}^{t},\tilde{y}_{i}^{t})-\nabla g_{i}^{t}(R_{i}\boldsymbol{x}_{i}^{t})\tilde{z}_{i}^{t}\big{)}\Big{\}}$,
10:     Send obscured parameters $y_{i}^{t+1}\!+\!\chi_{i}^{t+1}$, $z_{i}^{t+1}\!+\!\zeta_{i}^{t+1}$, and $\boldsymbol{x}_{i}^{t+1}\!+\boldsymbol{\vartheta}_{i}^{t+1}$ to its neighbors $j\in\mathcal{N}_{i}$.
11:end for
@@
In Algorithm 1, the matrix $R_{i}$ is used to extract the $i$th local variable $x_{i}^{t}$ from the variable $\boldsymbol{x}_{i}^{t}$, i.e., $R_{i}\boldsymbol{x}_{i}^{t}=x_{i}^{t}$. In this case, line 4 enables agent $i$ to track the weighted aggregative function $\frac{\lambda_{y}^{t}}{m}\sum_{i=1}^{m}g_{i}^{t}(x_{i}^{t})$ and line 5 allows agent $i$ to track the weighted average gradient $\frac{\lambda_{z}^{t}}{m}\sum_{i=1}^{m}\nabla_{y}f_{i}^{t}(x_{i}^{t},\tilde{y}_{i}^{t})$. It is clear that we use robust tracking approaches in line 4 and line 5 to update $y_{i}^{t}$ and $z_{i}^{t}$, respectively, which is different from existing distributed aggregative optimization algorithms [1, 5, 4, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 21, 22, 16] which use conventional gradient tracking techniques. This modification is necessary because the injected Laplace noise will accumulate over time in conventional gradient-tracking-based algorithms (see our prior works [47, 58] and others [59, 60] for a more detailed discussion on this issue). In fact, our experimental results in Fig. 1 and Fig. 2 have shown that existing distributed aggregative optimization algorithms in [4] and [21] (both of which rely on conventional gradient tracking techniques) exhibit divergent behaviors under Laplace noise. In addition, it is worth noting that since we use robust tracking approaches, our algorithmic structure (i.e., two robust tracking steps followed by a dynamic average-consensus gradient descent) is fundamentally different from that of existing distributed aggregative optimization algorithms in [1, 5, 4, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 21, 22, 16], which makes it impossible to use proof techniques for existing results.

### v3

@@ I Introduction
Distributed aggregative optimization addresses cooperative optimization and control problems where each agent’s objective function depends on both its local optimization variable and an aggregate of all agents’ optimization variables. Mathematically, the problem of distributed aggregative optimization can be formulated as follows:
$\min_{\boldsymbol{x}\in\Omega}~F(\boldsymbol{x})\!=\!\sum_{i=1}^{m}f_{i}(x_{i},g(\boldsymbol{x})),~~g(\boldsymbol{x})\!=\!\frac{1}{m}\sum_{i=1}^{m}g_{i}(x_{i}).$ (1)
Here, $m$ denotes the number of agents, $\boldsymbol{x}=\mathrm{col}(x_{1},\cdots,x_{m})$ denotes the concatenated optimization variable, where $x_{i}\in\Omega_{i}$ with $\Omega_{i}\subseteq\mathbb{R}^{n_{i}}$ denoting the constraint set. $g(\boldsymbol{x}):\Omega\mapsto\mathbb{R}^{r}$ is an aggregative function with $\Omega$ denoting the Cartesian product of all $\Omega_{i}$ (i.e., $\Omega=\Omega_{1}\times\cdots\times\Omega_{m}$) and each $g_{i}(x_{i}):\Omega_{i}\mapsto\mathbb{R}^{r}$ denoting the agent $i$’s private function. $f_{i}(x_{i},g(\boldsymbol{x})):\Omega_{i}\times\mathbb{R}^{r}\mapsto\mathbb{R}$ is the local objective function of agent $i$. Problem (1) has found various applications, with typical examples including warehouse placement optimization [35], multi-vehicle charging coordination [41, 13], cooperative robot surveillance [6], and personalized machine learning [8].
@@
To date, plenty of distributed aggregative optimization approaches have been proposed, for both time-invariant [35, 52, 7, 10, 56, 4, 18, 69, 68, 32, 23] and time-varying [6, 8, 63] objective functions. These approaches typically require participating agents to have access to accurate gradients of the objective functions, which, however, may not be feasible in practical applications. For example, in machine learning applications involving massive datasets, gradients are typically computed using randomly sampled data, which inherently introduces noise. Moreover, the use of mini-batch computations adds further randomness, increasing the variability in the gradient estimates available to the agents [5, 59, 33, 24]. These challenges necessitate a reevaluation of traditional approaches that assume accurate gradient information, thereby motivating the development of distributed stochastic aggregative optimization algorithms.
@@
However, developing an effective distributed stochastic algorithm for aggregative optimization is nontrivial due to the unique compositional structure of the objective functions. Specifically, unlike traditional distributed stochastic optimization [5, 59, 33, 24], the objective functions in distributed aggregative optimization explicitly depend on an inner-level aggregative function $g(\boldsymbol{x})$ (see (1)), which naturally arises in many practical applications. For example, in multi-vehicle charging coordination [41], each user’s cost (objective) function typically relies on electricity pricing, which usually varies as a function of the aggregated charging demand. This compositional dependence makes the gradient computation of objective functions inseparable from the estimation of the aggregative function. In the presence of noise, this inseparability leads to tangled errors in computed gradients and aggregative function estimates, substantially complicating the analysis. Moreover, in practice, computations must be conducted on sampled data, which introduces additional complexity. This is because at each iteration, both the aggregative function and the objective functions are defined by the sampled data, making them inherently time-varying and interdependent across iterations. Consequently, this creates significant challenges for algorithm design and convergence analysis. Although some progress has been made to address noise in distributed aggregative optimization [36, 11], these results only consider the special case where the objective function is stochastic and ignores noise in aggregative function estimation, which significantly simplifies the convergence analysis. In fact, when there is noise in aggregative function estimation, the existing approach in [36] will be subject to steady-state optimization errors. To the best of our knowledge, no results are available that can address stochasticity in both the objective functions and the aggregative function in distributed aggregative optimization.
@@
Notations: We use $\mathbb{R}^{n}$ to denote the Euclidean space of dimension $n$ and $\mathbb{N}$ ($\mathbb{N}^{+}$) to denote the set of non-negative (positive) integers. We use $\boldsymbol{1}_{m}$ and $\boldsymbol{0}_{m}$ to denote $m$-dimensional column vectors of all ones and all zeros, respectively. We let $W^{\top}$ denote the transpose of a matrix $W$. We use $\langle\cdot,\cdot\rangle$ to denote the inner product of two vectors and $\|\cdot\|$ to denote the Euclidean norm of a vector. We write $\text{col}(x_{1},\cdots,x_{m})$ for the stacked column vector of $x_{1},\cdots,x_{m}$. We denote $\mathbb{P}[\mathcal{A}]$ as the probability of an event $\mathcal{A}$ and $\mathbb{E}[x]$ as the expected value of a random variable $x$. $\Pi_{\Omega}$ denotes the Euclidean projection of a vector $x\in\mathbb{R}^{n}$ onto a compact set $\Omega$, i.e., $\Pi_{\Omega}:=\arg\min_{y\in\Omega}\|x-y\|$. We use $\text{Lap}(\nu_{i})$ to denote the Laplace distribution with a parameter $\nu_{i}>0$, featuring a probability density function $\frac{1}{2\nu_{i}}e^{\frac{-|x|}{\nu_{i}}}$. $\text{Lap}(\nu_{i})$ has a mean of zero and a variance of $2\nu_{i}^{2}$. We add an overbar to a letter to denote the average of $m$ agents, e.g., $\bar{y}^{t}=\frac{1}{m}\sum_{i=1}^{m}y_{i}^{t}$, and use bold font to represent stacked vectors of $m$ agents, e.g., $\boldsymbol{y}^{t}=\mathrm{col}(y_{1}^{t},\cdots,y_{m}^{t})$.
@@ II-A Distributed stochastic aggregative optimization
We consider distributed aggregative optimization problem in (1). In machine learning applications, the functions $f_{i}$ and $g_{i}$ are typically given by
@@
In Assumption 1, the Lipschitz continuity of the objective function is a desirable property in machine learning applications, as it often helps to improve the robustness and generalization of the learned model [61]. Various techniques have been developed to enforce this property, such as spectral normalization [48], spectral norm regularization [50], and gradient penalty methods [21]. Separately, the Lipschitz continuity of gradients is a standard assumption for convergence analysis in existing distributed aggregative optimization works (see, e.g., [35, 52, 7, 10, 56, 4, 18, 69, 68, 32, 23, 6, 8, 63, 36, 11]). Furthermore, when the constraint set $\Omega_{i}$ is compact, many loss functions and their corresponding gradients in machine learning satisfy this assumption, with typical examples including the RoBoSS loss function [1] and the widely used cross-entropy loss and its variants [3]. In addition, unlike existing distributed aggregative optimization results, in, e.g., [8, 6, 35, 10, 56, 18, 69, 32, 23, 52, 4, 63, 68, 36, 11] that require the objective functions to be strongly convex or convex, we allow both $f_{i}$ and $g_{i}$ to be nonconvex.
@@
The communication graph is undirected and connected. Moreover, the matrix $I+W$ is doubly stochastic and satisfies $\|I+W-\frac{\boldsymbol{1}_{m}\boldsymbol{1}_{m}^{\top}}{m}\|<1$.
According to Lemma 4 in [55], we have $\|I_{m}+W-\frac{\boldsymbol{1}_{m}\boldsymbol{1}_{m}^{\top}}{m}\|<1-|\rho_{2}|$, where $\rho_{2}$ is $W$’s second largest eigenvalue.
@@ II-B Local differential privacy (LDP)
$\vskip-1.99997pt\mathbb{P}[\mathcal{A}_{i}(\mathcal{D}_{i},\theta_{-i})\in\mathcal{O}_{i}]\leq e^{\epsilon_{i}}\mathbb{P}[\mathcal{A}_{i}(\mathcal{D}_{i}^{\prime},\theta_{-i})\in\mathcal{O}_{i}].$
@@ III Locally Differentially Private Distributed Aggregative Optimization Algorithm
We propose Algorithm 1 to solve the stochastic aggregative optimization problem (1) while ensuring rigorous $\epsilon_{i}$-LDP with a finite cumulative privacy budget in the infinite time horizon.
1:Input: Random initialization $\mathbf{x}_{i}^{0}\in\Omega$, $y_{i}^{0}\in{\mathbb{R}^{r}}$, and $z_{i}^{0}\in{\mathbb{R}^{r}}$; matrix $R_{i}\!=\![\boldsymbol{0}_{n_{i}\times n_{1}},\cdots,I_{n_{i}},\cdots,\boldsymbol{0}_{n_{i}\times n_{m}}]\!\in\!\mathbb{R}^{n_{i}\times n}$; stepsizes $\lambda_{x}^{t}=\frac{\lambda_{x}^{0}}{(t+1)^{v_{x}}}$, $\lambda_{y}^{t}=\frac{\lambda_{y}^{0}}{(t+1)^{v_{y}}}$, and $\lambda_{z}^{t}=\frac{\lambda_{z}^{0}}{(t+1)^{v_{z}}}$ with $\lambda_{x}^{0},\lambda_{y}^{0},\lambda_{z}^{0}>0$ and $v_{x},v_{y},v_{z}\in(0,1)$; DP noises $\chi_{i}^{t}\in\mathbb{R}^{r}$, $\zeta_{i}^{t}\in\mathbb{R}^{r}$, and $\boldsymbol{\vartheta}_{i}^{t}\in\mathbb{R}^{n}$ (with $n=\sum_{i=1}^{m}n_{i}$; here, we use bold font to represent stacked vectors of $m$ agents) satisfying Assumption 4.
2:for $t=0,1,\cdots,T-1$ do
4:  $y_{i}^{t+1}=y_{i}^{t}+\sum_{j\in\mathcal{N}_{i}}w_{ij}(y_{j}^{t}+\chi_{j}^{t}-y_{i}^{t})+\lambda_{y}^{t}g_{i}^{t}(R_{i}\mathbf{x}_{i}^{t})$;
6:     $\lambda_{z}^{t}\nabla_{y}f_{i}^{t}(R_{i}\mathbf{x}_{i}^{t},\tilde{y}_{i}^{t})$, with $\tilde{y}_{i}^{t}\triangleq\textstyle\frac{1}{\lambda_{y}^{t}}(y_{i}^{t+1}-y_{i}^{t})$;
7:  $\mathbf{x}_{i}^{t+1}=\Pi_{\Omega}\Big\{\mathbf{x}_{i}^{t}+\sum_{j\in\mathcal{N}_{i}}w_{ij}(\mathbf{x}_{j}^{t}+\boldsymbol{\vartheta}_{j}^{t}-\mathbf{x}_{i}^{t})$
8:     $\lambda_{x}^{t}R_{i}^{\top}\big(\nabla_{x}f_{i}^{t}(R_{i}\mathbf{x}_{i}^{t},\tilde{y}_{i}^{t})+\nabla g_{i}^{t}(R_{i}\mathbf{x}_{i}^{t})\tilde{z}_{i}^{t}\big)\Big\}$,
10:  Send obscured parameters $y_{i}^{t+1}\!+\!\chi_{i}^{t+1}$, $z_{i}^{t+1}\!+\!\zeta_{i}^{t+1}$, and $\mathbf{x}_{i}^{t+1}\!+\boldsymbol{\vartheta}_{i}^{t+1}$ to its neighbors $j\in\mathcal{N}_{i}$.
11:end for
The injected DP noises satisfy the following assumption:
@@
$\varsigma_{i,x}<v_{x}-v_{z},\quad\varsigma_{i,y}<v_{y},\quad\text{and}\quad\varsigma_{i,z}<v_{z},~\forall i\in[m],\vskip-3.00003pt$
@@
with $v_{x}$, $v_{y}$, and $v_{z}$ being the decaying rates of the stepsizes $\lambda_{x}^{t}$, $\lambda_{y}^{t}$ and $\lambda_{z}^{t}$, respectively, in Algorithm 1.
Remark 1.
The stepsizes in Algorithm 1, i.e., $\lambda_{x}^{t}$, $\lambda_{y}^{t}$, and $\lambda_{z}^{t}$, can be hard-coded offline in all agents’ programs, and hence, do not require any adjustment or coordination among agents during algorithmic implementation. Furthermore, since the stepsizes can be hard-coded prior to algorithmic implementation, heterogeneity in stepsizes across agents can be fully avoided and therefore does not to be considered in our design. In addition, we note that all noise parameters in Algorithm 1, i.e., $\sigma_{i,x}$, $\sigma_{i,y}$, $\sigma_{i,z}$, $\varsigma_{i,x}$, $\varsigma_{i,y}$, and $\varsigma_{i,z}$, can be locally selected by each agent $i$ based on its own practical privacy needs. Hence, our algorithm is inherently robust to heterogeneous noise parameters across agents.
@@
In line 6 of Algorithm 1, we use dynamic average-consensus-based gradient descent to enable each agent $i$ to estimate the optimization variables of all other agents, i.e., $\mathbf{x}_{i}^{t}\!=\!\mathrm{col}\{{\rm x}_{i(1)}^{t},\cdots,x_{i}^{t},\cdots,{\rm x}_{i(m)}^{t}\}$, where $x_{i}^{t}$ denotes agent $i$’s true optimization variable and ${\rm x}_{i(j)}^{t}$ denotes its estimate of agent $j$’s optimization variable for all $j\!\neq\!i$ and $j\!\in\![m]$. (Here, each agent $i$’s estimate of its own optimization variable coincides with the variable itself, i.e., ${\rm x}_{i(i)}^{t}\!=\!x_{i}^{t}$. Hence, we directly use $x_{i}^{t}$ to represent agent $i$’s estimate of its own optimization variable throughout the paper.) This estimation of neighbors’ optimization variables is an important novelty in our design (absent in all existing distributed aggregative optimization algorithms [35, 8, 6, 52, 7, 10, 56, 4, 18, 69, 68, 32, 23, 36, 11, 63]), and it is key to ensuring a finite cumulative privacy budget in the infinite time horizon. Specifically, by employing average consensus on $\mathbf{x}_{i}^{t}$, we can ensure a diminishing algorithmic sensitivity (see Eq. (16)), which, combined with meticulously designed diminishing DP-noise variances, can ensure a finite cumulative privacy budget in the infinite time horizon (see Eq. (21)). In contrast, in existing algorithms, the difference $\|x_{i}^{t}-x_{i}^{\prime t}\|$ of state trajectories under adjacent datasets will grow over time, which will in turn lead to a growing sensitivity, and hence, an exploding cumulative privacy budget as iterations proceed.
@@
In Algorithm 1, the matrix $R_{i}$ is used to extract the $i$th local variable $x_{i}^{t}$ from the variable $\mathbf{x}_{i}^{t}$, i.e., $R_{i}\mathbf{x}_{i}^{t}=x_{i}^{t}$. In this case, line 4 enables agent $i$ to track the weighted aggregative function $\frac{\lambda_{y}^{t}}{m}\sum_{i=1}^{m}g_{i}^{t}(x_{i}^{t})$ and line 5 allows agent $i$ to track the weighted average gradient $\frac{\lambda_{z}^{t}}{m}\sum_{i=1}^{m}\nabla_{y}f_{i}^{t}(x_{i}^{t},\tilde{y}_{i}^{t})$. It is clear that we use robust tracking approaches in line 4 and line 5 to update $y_{i}^{t}$ and $z_{i}^{t}$, respectively, which is different from existing distributed aggregative optimization algorithms [35, 8, 6, 52, 7, 10, 56, 4, 18, 69, 68, 32, 23, 36, 11, 63] which use conventional gradient tracking techniques. This modification is necessary because the injected Laplace noise will accumulate over time in conventional gradient-tracking-based algorithms. More specifically, we use the classical gradient-tracking-based distributed aggregative optimization algorithm in [36] as an example to illustrate the idea. In the absence of DP noises, the algorithm in [36] is given as follows:
$x_{i}^{t+1}$ $\textstyle=\Pi_{\Omega_{i}}\Big\{x_{i}^{t}-\alpha_{t}\left(\nabla_{x}f_{i}^{t}(x_{i}^{t},y_{i}^{t})+\nabla g_{i}(x_{i}^{t})z_{i}^{t}\right)\Big\},$
$y_{i}^{t+1}$ $\textstyle=\sum_{j=1}^{m}a_{ij}y_{j}^{t}+g_{i}(x_{i}^{t+1})-g_{i}(x_{i}^{t}),$
$z_{i}^{t+1}$ $\textstyle=\sum_{j=1}^{m}a_{ij}z_{j}^{t}+\nabla_{y}f_{i}^{t+1}(x_{i}^{t+1},y_{i}^{t+1})-\nabla_{y}f_{i}^{t}(x_{i}^{t},y_{i}^{t}),$
where $\sum_{j=1}^{m}a_{ij}=1$ and $\sum_{i=1}^{m}a_{ij}=1$ hold for all $i,j\in[m]$ according to Assumption 2 in [36] and the stepsize $\alpha_{t}$ is chosen as $\alpha_{t}=\frac{1}{\sqrt{t}}$ for all $t\geq 1$ following Eq. (8) in [36].
Using initialization $y_{i}^{0}\!=\!g_{i}(x_{i}^{0})$ and $z_{i}^{0}\!=\!\nabla_{y}f_{i}^{0}(x_{i}^{0},y_{i}^{0})$, we obtain $\bar{y}^{t}=\frac{1}{m}\sum_{i=1}^{m}g_{i}(x_{i}^{t})$ and $\bar{z}^{t}=\frac{1}{m}\sum_{i=1}^{m}\nabla_{y}f_{i}^{t}(x_{i}^{t},y_{i}^{t})$, which means that ensuring the consensus of $y_{i}^{t}$ and $z_{i}^{t}$, i.e., $y_{i}^{t}=\bar{y}^{t}$ and $z_{i}^{t}=\bar{z}^{t}$, is sufficient for each agent to accurately track the aggregative function and the global gradient, respectively, i.e., achieving $y_{i}^{t}=\bar{y}^{t}=\frac{1}{m}\sum_{i=1}^{m}g_{i}(x_{i}^{t})$ and $z_{i}^{t}=\bar{z}^{t}=\frac{1}{m}\sum_{i=1}^{m}\nabla_{y}f_{i}^{t}(x_{i}^{t},y_{i}^{t})$, respectively.
When noises are added to both shared variables $y_{i}^{t}$ and $z_{i}^{t}$ to ensure $\epsilon_{i}$-LDP, the algorithm in [36] becomes
$x_{i}^{t+1}$ $\textstyle\!\!=\!\Pi_{\Omega_{i}}\Big\{x_{i}^{t}-\alpha_{t}\left(\nabla_{x}f_{i}^{t}(x_{i}^{t},y_{i}^{t})+\nabla g_{i}(x_{i}^{t})z_{i}^{t}\right)\Big\},$
$y_{i}^{t+1}$ $\textstyle\!\!=\!\sum_{j=1}^{m}a_{ij}(y_{j}^{t}+\chi_{j}^{t})+g_{i}(x_{i}^{t+1})-g_{i}(x_{i}^{t}),$
$z_{i}^{t+1}$ $\textstyle\!\!=\!\sum_{j=1}^{m}a_{ij}(z_{j}^{t}\!+\!\zeta_{j}^{t})\!+\!\nabla_{y}f_{i}^{t+1}(x_{i}^{t+1}\!,y_{i}^{t+1})\!-\!\nabla_{y}f_{i}^{t}(x_{i}^{t},y_{i}^{t}),$
where $\chi_{j}^{t}\in\mathbb{R}^{r}$ and $\zeta_{j}^{t}\in\mathbb{R}^{r}$ represent the DP noises injected on variables $y_{j}^{t}$ and $z_{j}^{t}$, respectively.
It can be seen that under the same initialization $y_{i}^{0}=g_{i}(x_{i}^{0})$ and $z_{i}^{0}=\nabla_{y}f_{i}^{0}(x_{i}^{0},y_{i}^{0})$, the following relations hold true by induction:
$\bar{y}^{t}$ $\textstyle=\frac{1}{m}\sum_{i=1}^{m}\left(g_{i}(x_{i}^{t})+\sum_{k=0}^{t-1}\chi_{i}^{k}\right);$ (4)
$\bar{z}^{t}$ $\textstyle=\frac{1}{m}\sum_{i=1}^{m}\left(\nabla_{y}f_{i}^{t}(x_{i}^{t},y_{i}^{t})+\sum_{k=0}^{t-1}\zeta_{i}^{k}\right),$
which implies that the DP noises accumulate over time in the estimates of both the aggregative function and the global gradient. Therefore, when the variables $y_{i}^{t}$ and $z_{i}^{t}$ are directly fed into the update of the model parameter $x_{i}^{t+1}$, learning accuracy will be compromised. In fact, our experimental results in Fig. 1(a) and Fig. 2(a) have shown that existing distributed aggregative optimization algorithms in [6] and [36] (both of which rely on conventional gradient tracking techniques) exhibit divergent behaviors under Laplace DP noise.
In contrast, in Algorithm 1, we use the increments $y_{i}^{t+1}-y_{i}^{t}$ and $z_{i}^{t+1}-z_{i}^{t}$ in the update of the model parameter $x_{i}^{t+1}$. Because the two increments can be verified to satisfy (by using Line 4 and Line 5 of Algorithm 1 and Assumption 3):
$\bar{y}^{t+1}-\bar{y}^{t}$ $\textstyle=\frac{\lambda_{y}^{t}}{m}\sum_{i=1}^{m}\left(g_{i}^{t}(R_{i}\mathbf{x}_{i}^{t})+\sum_{j\in\mathcal{N}_{i}}w_{ij}\chi_{j}^{t}\right);$
$\bar{z}^{t+1}-\bar{z}^{t}$ $\textstyle=\frac{\lambda_{z}^{t}}{m}\sum_{i=1}^{m}\left(\nabla_{y}f_{i}^{t}(R_{i}\mathbf{x}_{i}^{t},\tilde{y}_{i}^{t})+\sum_{j\in\mathcal{N}_{i}}w_{ij}\zeta_{j}^{t}\right),$
this design effectively circumvents the noise accumulation problem discussed above in the existing algorithm of [36]. In addition, it is worth noting that since we use robust tracking approaches, our algorithmic structure (i.e., two robust tracking steps followed by a dynamic average-consensus gradient descent) is fundamentally different from that of existing distributed aggregative optimization algorithms in [35, 8, 6, 52, 7, 10, 56, 4, 18, 69, 68, 32, 23, 36, 11, 63], which makes it impossible to use proof techniques for existing results.